OpenStudy (anonymous):

integrate ((x-3)/(x-4))^(1/2)

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{((x-3)/(x-4))} dx\]

OpenStudy (anonymous):

(x-3)/(x-4) = 1 + 1/(x-4) let 1/(x-4)= tan^2y then, dx = 2tany sec^2y .dy

OpenStudy (anonymous):

SO,